# Maintenance Optimization for Asset Networks with Unknown Degradation Parameters

### v2

@@ 3.1 Compound Poisson degradation
We consider a set of assets (machines) $\mathcal{M}=\{1,\ldots,M\}$, each with a single critical component that degrades independently as random shocks arrive. From this point forward, we will refer to the degradation of components as asset degradation. Shocks occur according to a Poisson process, and the damage that accumulates during each shock is a non-negative random variable, meaning that degradation follows a compound Poisson processes. Such shock degradation is commonly modeled using Poisson processes in the literature; see, e.g., Tian et al. (2021), Drent et al. (2023), Drent et al. (2024b, a), and Wang et al. (2024). The Poisson intensity of shock arrivals at machine $m\in\mathcal{M}$ is denoted by $\lambda_{m}\in\mathbb{R}_{+}$. The random damage amount at machine $m$ resulting from a shock adheres to the distribution of a member of the one-parameter (denoted as $\phi_{m}$) exponential family. The probability density or mass function of such a random variable can be written as
@@
where $T_{m}(x)$ represents the sufficient statistic, and $h_{m}(x)$ and $A_{m}(\phi_{m})$ are known functions. We assume that shock sizes are non-negative and that the sufficient statistic is linear, i.e., $T_{m}(x)\equiv x$, which enables a state space reduction in our optimization problem. The function $T_{m}(\cdot)$ contains all the information necessary to compute any estimate of the parameter $\phi_{m}$. In the literature, this group of distributions is commonly known as the linear exponential family or natural exponential family, named for its linear sufficient statistic, and was first introduced by Morris (1982). This class includes several well-known distributions used in maintenance. To illustrate our approach, we consider the geometric distribution (supported on $\mathbb{Z}_{+}$) as a representative example. It is important to note that our work is not restricted to this particular distribution.
@@
$f(x\mid p_{m})=(1-p_{m})^{x}p_{m}=e^{\textrm{ln}(1-p_{m})x-ln(1/p_{m})}.$
Note that $h_{m}(x)=1$, $T_{m}(x)=x$, $\phi_{m}=\textrm{ln}(1-p_{m})$ and $A_{m}(\phi_{m})=\textrm{ln}(1/p_{m})$.
We assume continuous remote access to the degradation level for each machine through sensory equipment, however, interaction with the system is limited to evenly spaced decision epochs that correspond to scheduled maintenance activities. Without loss of generality, we rescale time such that the time between two consecutive decision epochs equals one. Let $N_{m}(t)$ denote the total number of shocks incurred by the $m$-th asset since its installation up to its operational age (that is, the time since the last replacement) $t\in\mathbb{R}_{+}$, i.e., $N_{m}(t)$ is a Poisson process with rate $\lambda_{m}$. Here, it is important to note that real-time continuous monitoring of the machine, rather than relying solely on periodic inspections, enables an accurate count of the total number of shocks. The $m$-th asset’s degradation level is denoted by $X_{m}(t)\in[0,\xi_{m}]$ for some failure threshold $\xi_{m}>0$. When the asset’s degradation level reaches or exceeds $\xi_{m}$, it breaks down and requires corrective maintenance at the next decision epoch. Maintenance is instantaneous and restores the asset to an as-good-as-new condition. Note that the assumptions of limited interaction with the system and instantaneous component replacements are reasonably applicable to certain asset networks, e.g., offshore wind farms. In the case study presented in Section 6, a time unit is defined to roughly correspond to the minimum time required to dispatch a maintenance team or deliver a spare part.
@@
where $X_{m}(0)=N_{m}(0)=0$ by definition. Furthermore, let $\bm{Y}_{m}\big{(}(t-1,t]\big{)}=(Y^{(N_{t-1}+1)}_{m},Y^{(N_{t-1}+2)}_{m},\ldots,Y^{(N_{t})}_{m})$ be the sizes of the shocks that arrive between age t1t-1 and $t$, and let $Z_{m}\big{(}(t-1,t]\big{)}=\sum_{i=N_{m}(t-1)+1}^{N_{m}(t)}Y^{(i)}_{m}$ be the corresponding total incurred damage. Let $k_{m}(t)$ denote the observed number of shocks of the $m$-th asset, that is, $k_{m}(t)$ is the realization of $K_{m}\big{(}(t-1,t]\big{)}$. Denote with $\bm{y}_{m}(t)=(y_{m}^{(1)},\ldots,y_{m}^{(k_{m}(t))})$ the corresponding observed shock sizes. That is, $\bm{y}_{m}(t)$ is the realization of $\bm{Y}_{m}\big{(}(t-1,t]\big{)}$. The tuple $\bm{\theta}_{m}(t)=(k_{m}(t),\bm{y_{m}}(t))$ is the observed degradation signal of machine $m$ between ages t1t-1 and $t$. Given the assumption of a linear sufficient statistic, we do not need to keep track of individual shock sizes. Instead, we can collapse the state space by summarizing the signal as $(k_{m}(t),z_{m}(t))$, where $z_{m}(t)=\sum_{i=1}^{k_{m}(t)}y_{m}^{(i)}(t)$ represents the accumulated damage between ages t1t-1 and $t$. That is, $z_{m}(t)$ is the realization of $Z_{m}\big{(}(t-1,t]\big{)}$. This state space collapse is a key simplification that enables more tractable inference.
Recall that each asset stems from a distinct heterogeneous population that consists of assets with different degradation parameters $\lambda_{m}$ and $\phi_{m}$. In practice, these parameters are hidden; only the observed degradation signal $\bm{\theta}_{m}(t)$ is available at operational age $t$. In this case, we will employ a POMDP to model the integrated challenge of learning the degradation parameters while determining the optimal timing for replacement. In the underlying MDP, these parameters $\lambda_{m}$ and $\phi_{m}$ are drawn from known distributions, $\Lambda_{m}$ and $\Phi_{m}$, respectively, and are observed by an oracle that is aware of the true population heterogeneity. The asset manager’s knowledge about the degradation model and its parameters gives rise to various information levels which are detailed in the next section.
@@ 3.2 Information levels
Inspired by the approach of Da Costa et al. (2023), we formalize the asset manager’s knowledge of the degradation model and its underlying parameters using three distinct levels of information:
@@
    The asset manager has no information about the model’s parameters $\lambda_{m},\phi_{m}$, nor any knowledge on their distribution (i.e., the distributions $\Lambda_{m}$ and $\Phi_{m}$ are unknown) for each $m\in\mathcal{M}$.
@@
    The asset manager has no information about the model’s parameters $\lambda_{m},\phi_{m}$, but has full knowledge on their distribution (i.e., the distributions $\Lambda_{m}$ and $\Phi_{m}$ are known) for each $m\in\mathcal{M}$.
@@
Given an information level $\textbf{L}\in\{\textbf{L}_{0},\textbf{L}_{1},\textbf{L}_{2}\}$, the objective is to devise a policy $\pi^{\textbf{L}}$ that minimizes the total expected discounted cost of managing a specific network of assets. The baseline information level, $\textbf{L}_{0}$, reflects the typical conditions encountered by the asset manager in practice. To address the heterogeneity of the asset population, hyperparameters of the distributions $\Lambda_{m}$ and $\Phi_{m}$ for each $m\in\mathcal{M}$ should be estimated from available historical degradation data. Subsequently, the asset manager formulates and solves an approximate problem under the enhanced information level $\textbf{L}_{1}$, and implements the resulting solution in practice. In contrast, information level $\textbf{L}_{2}$ provides the most accurate and informed basis for decision-making. Consequently, an optimal policy derived under this level of information achieves the lowest total expected discounted cost. This progression from $\textbf{L}_{0}$ to $\textbf{L}_{2}$ highlights the trade-off between the costs of acquiring additional information and the benefits of an improved CBM planning.
@@ 3.3 Partially observable Markov decision process formulation
The state-dependent action set $\mathcal{U}(h)=\mathcal{U}_{1}(h)\times\ldots\times\mathcal{U}_{M}(h)$ is formed by taking the Cartesian product of the $M$ individual state-dependent action sets.
The asset manager incurs costs related to either the corrective or preventive replacement of an asset. If the degradation level at a decision epoch is less than $\xi_{m}$, then we can either maintain the asset preventively at cost $c_{m}^{\textrm{PM}}$ or proceed to the next decision epoch without incurring any cost. If the degradation level of machine $m$ at a decision epoch is greater than or equal to the failure threshold $\xi_{m}$, then the failed asset is replaced correctively at cost $c_{m}^{\textrm{CM}}$. We assume that for all $m\in\mathcal{M}$ it holds that $0<c_{m}^{\textrm{PM}}<c_{m}^{\textrm{CM}}<\infty$ to avoid unrealistic cases. Moreover, both types of replacements take negligible time. The assets are economically coupled through a shared setup cost, a widely studied form of economic dependence in the literature. Specifically, if at least one asset is replaced during a decision epoch, a one-time setup cost $c^{\textrm{ST}}\geq 0$ is incurred. Although we focus on this particular economic dependence, the algorithmic developments presented in this paper readily extend to other forms of economic dependencies, such as more complex setup cost functions. After an asset $m$ is replaced, new degradation parameters $\lambda_{m}$ and $\phi_{m}$ are drawn from their respective distributions, $\Lambda_{m}$ and $\Phi_{m}$. Therefore, the costs incurred when taking action $a=(a_{1},\ldots,a_{M})\in\mathcal{U}(h)$ in state $h$ are:
@@ 3.4 Bayesian Markov decision process formulation
The use of conjugate pairs refines this integration by ensuring that the updated belief distributions remain within the same family as the prior belief distributions. For example, the gamma distribution acts as a conjugate prior for the (unknown) rate parameter of a Poisson distribution. Thus, for each asset $m\in\mathcal{M}$, we assume that $\lambda_{m}$ is drawn from a gamma distribution, $\Lambda_{m}\sim\textrm{Gamma}(\alpha_{m},\beta_{m})$, where $\alpha_{m}>0$ is the shape parameter and $\beta_{m}>0$ is the scale parameter. Additionally, for each $m\in\mathcal{M}$, we assume that $\Phi_{m}$ is distributed according to the general prior for a member of the exponential family characterized by hyperparameters $a_{m}>0$ and $b_{m}>0$. A member of the exponential family has a conjugate prior with a density that can be expressed as
$f_{\Phi_{m}}\left(\phi_{m}\mid a_{m}(t),b_{m}(t)\right)=H_{m}\left(a_{m}(t),b_{m}(t)\right)e^{a_{m}(t)\phi_{m}-b_{m}(t)A_{m}(\phi_{m})},$
where $H_{m}\left(a_{m}(t),b_{m}(t)\right)$ is a normalizing constant (Ghosh et al., 2007). The beta distribution serves as a conjugate prior for the parameter of several distributions in the exponential family, for instance the geometric distribution.
When a new asset $m$ is installed, the parameters $\lambda_{m}$ and $\phi_{m}$ of the compound Poisson degradation process are drawn from the known distributions $\Lambda_{m}$ and $\Phi_{m}$, and there is no available history. Thus, we interpret the parameters $\lambda_{m}$ and $\phi_{m}$ as random variables, denoted by $\tilde{\Lambda}_{m}$ and $\tilde{\Phi}_{m}$.
@@
At time $t=0$, the manager has a prior belief about the degradation process parameters of each asset $m$, which is expressed as a probability distribution over all possible combinations. The joint prior belief distribution of asset $m$ satisfies
$f^{0}_{\tilde{\Lambda}_{m},\tilde{\Phi}_{m}}(\lambda_{m},\phi_{m}):=f_{\Lambda_{m}}(\lambda_{m}\mid\alpha_{m},\beta_{m})\cdot f_{\Phi_{m}}(\phi_{m}\mid a_{m},b_{m}).$
@@
At time $t$, we use the observed degradation signal $\bm{\theta}_{m}(t)$ to infer the joint distribution of $\tilde{\Lambda}_{m}$ and $\tilde{\Phi}_{m}$. As shown by Drent et al. (2023, Proposition 1), this joint distribution can be factorized into two independent distributions of the same form, with parameters updated solely based on the information contained in $\bm{\theta}_{m}(t)$ observed in the previous time period.
@@
$f^{t}_{\tilde{\Lambda}_{m},\tilde{\Phi}_{m}}(\lambda_{m},\phi_{m})=f^{t-1}_{\tilde{\Lambda}_{m}}(\lambda_{m}\mid\alpha_{m}+k_{m}(t),\beta_{m}+t_{m}(t))\cdot f^{t-1}_{\tilde{\Phi}_{m}}(\phi_{m}\mid a_{m}+x_{m}(t),b_{m}+k_{m}(t)).$ (3)
@@
$b_{m}(t)=b_{m}+k_{m}(t).$ (7)
Thus, the state $\tilde{h}(t)$ at time $t$ of the associated BMDP can be represented by a vector $\tilde{h}(t)=(x_{1}(t),k_{1}(t),t_{1}(t),\ldots,x_{M}(t),k_{M}(t),t_{M}(t))$. The objective of Eq. (2) is equivalent to

### v3

@@ 3.1 Compound Poisson degradation
We consider a set of assets (machines) $\mathcal{M}=\{1,\ldots,M\}$, each with a single critical component that degrades independently as random shocks arrive. From this point forward, we will refer to the degradation of components as asset degradation. Shocks occur according to a Poisson process, and the damage that accumulates during each shock is a non-negative random variable, meaning that degradation follows a compound Poisson process. This type of shock model is appropriate for, e.g., certain metal and ceramic components in trains, aircraft, and medical equipment (including the IXR systems of our case study) that primarily deteriorate when subjected to discrete stress events [Esary and Marshall, 1973]. In such settings, degradation accumulates through a sequence of shocks associated with random events, such as thermal cycles, mechanical loads, or electrical stress. Modeling these shock arrivals as a Poisson process is common in the reliability and maintenance literature, as it provides a tractable representation of random and approximately independent stress events occurring over time. Indeed, shock-based degradation models with Poisson arrivals have been widely used in maintenance optimization; see, e.g., [Tian et al., 2021, Drent et al., 2023, 2024a, 2024b, Wang et al., 2024]. Consistent with this literature, we adopt the same modeling approach. The Poisson intensity of shock arrivals at machine $m\in\mathcal{M}$ is denoted by $\lambda_{m}\in\mathbb{R}_{+}$. The random damage amount at machine $m$ resulting from a shock adheres to the distribution of a member of the one-parameter (denoted as $\phi_{m}$) exponential family. The probability density or mass function of such a random variable can be written as
@@
where $T_{m}(x)$ represents the sufficient statistic, and $h_{m}(x)$ and $A_{m}(\phi_{m})$ are known functions. We assume that shock sizes are non-negative and that the sufficient statistic is linear, i.e., $T_{m}(x)\equiv x$, which enables a state space reduction in our optimization problem. The function $T_{m}(\cdot)$ contains all the information necessary to compute any estimate of the parameter $\phi_{m}$. In the literature, this group of distributions is commonly known as the linear exponential family or natural exponential family, named for its linear sufficient statistic, and was first introduced by Morris [1982]. To illustrate our approach, we consider the geometric distribution (supported on $\mathbb{Z}_{+}$) as a representative example.
@@
$f(x\mid p_{m})=(1-p_{m})^{x}p_{m}=e^{\ln(1-p_{m})x-\ln(1/p_{m})}.$
Note that $h_{m}(x)=1$, $T_{m}(x)=x$, $\phi_{m}=\ln(1-p_{m})$ and $A_{m}(\phi_{m})=\ln(1/p_{m})$.
Remark
The geometric distribution serves as an illustrative example, but the framework extends to any one-parameter exponential family with linear sufficient statistic $T_{m}(x)=x$. This class includes several commonly used discrete and continuous distributions for modeling degradation increments. In the discrete case, examples of the linear exponential family include the Poisson distribution, the binomial distribution (with fixed number of trials), and the negative binomial distribution (with fixed number of successes). In the continuous case, the exponential distribution, normal distribution (with fixed variance), inverse Gaussian (with fixed shape parameter), and the gamma distribution (with fixed shape parameter) are all members of the linear exponential family. In C, we demonstrate for two standard discrete distributions—the binomial and the Poisson—how they can be written in canonical exponential-family form with linear sufficient statistic, thereby fitting within our modeling framework.
We assume continuous, perfect, remote access to the degradation level for each machine through sensory equipment, while interaction with the system is limited to evenly spaced decision epochs (without loss of generality, we rescale time such that the time between two consecutive decision epochs equals one) corresponding to scheduled maintenance opportunities. Two remarks are in order to motivate this assumption. First, regular maintenance intervals are common in practice when maintenance planning and logistics must be arranged well in advance. For example, in geographically dispersed asset networks, maintenance visits are often scheduled at fixed intervals (e.g., semi-annually), as mobilizing crews and spare parts requires substantial preparation. Second, with the increasing deployment of internet-connected assets, embedded sensors, and digital monitoring infrastructures, near-continuous and accurate condition monitoring is becoming standard in many industrial applications. We therefore deliberately study this data-rich setting as a benchmark for integrated learning and decision-making. In the case study presented in Section 6, a time unit is defined to roughly correspond to the minimum time required to dispatch a maintenance team or deliver a spare part.
Let $N_{m}(t)$ denote the total number of shocks incurred by the $m$-th asset since its installation up to its operational age (that is, the time since the last replacement) $t\in\mathbb{R}_{+}$, i.e., $N_{m}(t)$ is a Poisson process with rate $\lambda_{m}$. Here, it is important to note that real-time continuous monitoring of the machine, rather than relying solely on periodic inspections, enables an accurate count of the total number of shocks. The $m$-th asset’s degradation level is denoted by $X_{m}(t)\in[0,\xi_{m}]$ for some failure threshold $\xi_{m}>0$. When the asset’s degradation level reaches or exceeds $\xi_{m}$, it breaks down and requires corrective maintenance at the next decision epoch. Maintenance is instantaneous and restores the asset to an as-good-as-new condition.
@@
where $X_{m}(0)=0$ and $N_{m}(0)=0$ by definition. Furthermore, let $\bm{Y}_{m}\big((t-1,t]\big)=(Y^{(N_{m}(t-1)+1)}_{m},\ldots,Y^{(N_{m}(t))}_{m})$ be the sizes of the shocks that arrive between age t1t-1 and $t$, and let $Z_{m}\big((t-1,t]\big)=\sum_{i=N_{m}(t-1)+1}^{N_{m}(t)}Y^{(i)}_{m}$ be the corresponding total incurred damage. Let $k_{m}(t)$ denote the observed number of shocks of the $m$-th asset, that is, $k_{m}(t)$ is the realization of $K_{m}\big((t-1,t]\big)$. Denote with $\bm{y}_{m}(t)=(y_{m}^{(1)},\ldots,y_{m}^{(k_{m}(t))})$ the corresponding observed shock sizes. That is, $\bm{y}_{m}(t)$ is the realization of $\bm{Y}_{m}\big((t-1,t]\big)$. The tuple $\bm{\theta}_{m}(t)=(k_{m}(t),\bm{y}_{m}(t))$ is the observed degradation signal of machine $m$ between ages t1t-1 and $t$. Given the assumption of a linear sufficient statistic, we do not need to keep track of individual shock sizes. Instead, we can collapse the state space by summarizing the signal as $(k_{m}(t),z_{m}(t))$, where $z_{m}(t)=\sum_{i=1}^{k_{m}(t)}y_{m}^{(i)}(t)$ represents the accumulated damage between ages t1t-1 and $t$. That is, $z_{m}(t)$ is the realization of $Z_{m}\big((t-1,t]\big)$. This state space collapse is a key simplification that enables more tractable inference.
Recall that each component stems from a distinct heterogeneous population that consists of components with different degradation parameters $\lambda_{m}$ and $\phi_{m}$. In practice, these parameters are hidden; only the observed degradation signal $\bm{\theta}_{m}(t)$ is available at operational age $t$. In this case, we will employ a POMDP to model the integrated challenge of learning the degradation parameters while determining the optimal timing for replacement. In the underlying MDP, these parameters $\lambda_{m}$ and $\phi_{m}$ are drawn from known distributions, $\Lambda_{m}$ and $\Phi_{m}$, respectively, and are observed by an oracle that is aware of the true population heterogeneity. The asset manager’s knowledge about the degradation model and its parameters gives rise to various information levels, which are detailed in the next section.
@@ 3.2 Information levels
Inspired by the approach of Da Costa et al. [2023], we formalize the asset manager’s knowledge of the degradation model and its underlying parameters using three distinct levels of information:
@@
    The asset manager has no information about the model’s parameters $\lambda_{m},\phi_{m}$, nor any knowledge of their distribution (i.e., the distributions $\Lambda_{m}$ and $\Phi_{m}$ are unknown) for each $m\in\mathcal{M}$.
@@
    The asset manager has no information about the model’s parameters $\lambda_{m},\phi_{m}$, but has full knowledge of their distribution (i.e., the distributions $\Lambda_{m}$ and $\Phi_{m}$ are known) for each $m\in\mathcal{M}$.
@@
Given an information level $\textbf{L}\in\{\textbf{L}_{0},\textbf{L}_{1},\textbf{L}_{2}\}$, the objective is to devise a policy $\pi^{\textbf{L}}$ that minimizes the total expected discounted cost of managing a specific network of assets. The baseline information level, $\textbf{L}_{0}$, reflects the typical conditions encountered by the asset manager in practice. To address the heterogeneity of the component population, hyperparameters of the distributions $\Lambda_{m}$ and $\Phi_{m}$ for each $m\in\mathcal{M}$ should be estimated from available historical degradation data. Subsequently, the asset manager formulates and solves an approximate problem under the enhanced information level $\textbf{L}_{1}$, and implements the resulting solution in practice. In contrast, information level $\textbf{L}_{2}$ provides the most accurate and informed basis for decision-making. Consequently, an optimal policy derived under this level of information achieves the lowest total expected discounted cost. This progression from $\textbf{L}_{0}$ to $\textbf{L}_{2}$ highlights the trade-off between the costs of acquiring additional information and the benefits of improved CBM planning. However, although relevant, these acquisition costs are not explicitly incorporated in our model and are left for future work.
@@ 3.3 Partially observable Markov decision process formulation
The state-dependent action set $\mathcal{U}(h)=\mathcal{U}_{1}(h)\times\ldots\times\mathcal{U}_{M}(h)$ is formed by taking the Cartesian product of the $M$ individual state-dependent action sets. For tractability, we assume that maintenance capacity is sufficient to service all assets simultaneously.
The asset manager incurs costs related to either the corrective or preventive replacement of an asset. If the degradation level at a decision epoch is less than $\xi_{m}$, then we can either maintain the asset preventively at cost $c_{m}^{\textrm{PM}}$ or proceed to the next decision epoch without incurring any cost. If the degradation level of machine $m$ at a decision epoch is greater than or equal to the failure threshold $\xi_{m}$, then the failed asset is replaced correctively at cost $c_{m}^{\textrm{CM}}$. We assume that for all $m\in\mathcal{M}$ it holds that $0<c_{m}^{\textrm{PM}}<c_{m}^{\textrm{CM}}<\infty$ to avoid unrealistic cases. This cost structure is commonly adopted in the maintenance literature [De Jonge and Scarf, 2020]. Moreover, both types of replacements take negligible time, which is a reasonable assumption because in practice, replacement times are relatively small compared to the time between decision epochs. The assets are economically coupled through a shared setup cost, a widely adopted assumption for modeling economic dependence in multi-component maintenance systems [Olde Keizer et al., 2017]. In practice, such setup costs represent fixed expenses associated with initiating maintenance activities, such as dispatching a maintenance crew, preparing equipment, or temporarily shutting down part of the system. Specifically, if at least one asset is replaced during a decision epoch, a one-time setup cost $c^{\textrm{ST}}\geq 0$ is incurred. Although we focus on this particular economic dependence, the algorithmic developments presented in this paper readily extend to other forms of economic dependencies, such as more complex setup cost functions. After an asset $m$ is replaced, new degradation parameters $\lambda_{m}$ and $\phi_{m}$ are drawn from their respective distributions, $\Lambda_{m}$ and $\Phi_{m}$. Therefore, the costs incurred when taking action $a=(a_{1},\ldots,a_{M})\in\mathcal{U}(h)$ in state $h$ are:
@@ 3.4 Bayesian Markov decision process formulation
The use of conjugate pairs refines this integration by ensuring that the updated belief distributions remain within the same family as the prior belief distributions. For example, the gamma distribution acts as a conjugate prior for the (unknown) rate parameter of a Poisson distribution. Thus, for each asset $m\in\mathcal{M}$, we assume that $\lambda_{m}$ is drawn from a gamma distribution, $\Lambda_{m}\sim\textrm{Gamma}(\alpha_{m},\beta_{m})$, where $\alpha_{m}>0$ is the shape parameter and $\beta_{m}>0$ is the scale parameter. Additionally, for each $m\in\mathcal{M}$, we assume that $\Phi_{m}$ is distributed according to the general prior for a member of the exponential family characterized by hyperparameters $r_{m}>0$ and $s_{m}>0$. A member of the exponential family has a conjugate prior with a density that can be expressed as
$f_{\Phi_{m}}\left(\phi_{m}\mid r_{m}(t),s_{m}(t)\right)=H_{m}\left(r_{m}(t),s_{m}(t)\right)e^{r_{m}(t)\phi_{m}-s_{m}(t)A_{m}(\phi_{m})},$
where $H_{m}\left(r_{m}(t),s_{m}(t)\right)$ is a normalizing constant [Ghosh et al., 2007]. The beta distribution serves as a conjugate prior for the parameter of several distributions in the exponential family, for instance the geometric distribution.
When asset $m$ is installed, the parameters $\lambda_{m}$ and $\phi_{m}$ of the compound Poisson degradation process are drawn from the known distributions $\Lambda_{m}$ and $\Phi_{m}$, and there is no available history. Thus, we interpret the parameters $\lambda_{m}$ and $\phi_{m}$ as random variables, denoted by $\tilde{\Lambda}_{m}$ and $\tilde{\Phi}_{m}$.
@@
At time $t=0$, the manager has a prior belief about the degradation process parameters of each asset, which is expressed as a probability distribution over all possible combinations. The joint prior belief distribution of asset $m$ satisfies
$f^{(0)}_{\tilde{\Lambda}_{m},\tilde{\Phi}_{m}}(\lambda_{m},\phi_{m}):=f_{\Lambda_{m}}(\lambda_{m}\mid\alpha_{m},\beta_{m})\cdot f_{\Phi_{m}}(\phi_{m}\mid r_{m},s_{m}).$
@@
At time $t$, we use the observed degradation signal $\bm{\theta}_{m}(t)$ to infer the joint distribution of $\tilde{\Lambda}_{m}$ and $\tilde{\Phi}_{m}$. As shown by Drent et al. [2023, Proposition 1], this joint distribution can be factorized into two independent distributions of the same form, with parameters updated solely based on the information contained in $\bm{\theta}_{m}(t)$ observed in the previous time period.
@@
$f^{(t)}_{\tilde{\Lambda}_{m},\tilde{\Phi}_{m}}(\lambda_{m},\phi_{m})=f^{(t-1)}_{\tilde{\Lambda}_{m}}(\lambda_{m}\mid\alpha_{m}+k_{m}(t),\beta_{m}+t_{m}(t))\cdot f^{(t-1)}_{\tilde{\Phi}_{m}}(\phi_{m}\mid r_{m}+x_{m}(t),s_{m}+k_{m}(t)).$ (3)
@@
$s_{m}(t)=s_{m}+k_{m}(t).$ (7)
Thus, the state at time $t$ of the associated BMDP can be represented by a vector $\tilde{h}(t)=(x_{1}(t),k_{1}(t),t_{1}(t),\ldots,x_{M}(t),k_{M}(t),t_{M}(t))$. The objective of Eq. (2) is equivalent to